# Towards Autonomous Driving with Small-Scale Cars: A Survey of Recent Development

### v1

@@ III-E Obstacle Detection and Avoidance
III-E1 Vision-based approach
@@
The vision-based approach relies on images captured by cameras as the primary input for detection. This approach employs computer vision (CV) or ML algorithms, prominently leveraging CNN, for effective detection tasks. Within the realm of cameras used, the vision-based approach can be categorized into two subclasses: monocular and stereo. Monocular image-based methods rely on a single image for processing, while stereo methods utilize images captured by two synchronized cameras. Details about the camera sensors are discussed in Section IV-A.
@@
Within monocular image-based methods, as outlined in [191], the obstacle detection task for small-scale cars is typically categorized into three primary domains: appearance-based, motion-based, and depth-based methods. With appearance-based methods, obstacles are typically identified as the foreground within images. The primary challenge lies in distinguishing relevant foreground or background elements based on established criteria, such as color discrepancies [192] or texture features [193]. In the work by [36], the RGB images captured by the monocular camera of a Duckiebot are initially transformed into the HSV colorspace. Subsequently, the color filter from OpenCV is employed to detect obstacles based on color differences. However, the drawback is evident: success detection depends on the predefined color of the obstacles, limiting adaptability to diverse scenarios. Motion-based obstacle detection refers to identifying obstacles or objects in an environment by analyzing motion vectors in the image, it involves comparing successive frames of images to determine alterations in position, velocity, or other motion-related characteristics. Objects that move or exhibit changes in their motion characteristics are then identified as potential obstacles. Optical flow analysis [194], background subtraction, differential methods. can be employed for motion-based obstacle detection [195, 196]. Depth-based methods utilize depth information extracted from images to discern the distances and spatial arrangement of objects within the environment, aiding in accurate object detection [197, 198]. Employing conventional image processing techniques like this may fall short in meeting real-time application expectations, primarily due to their inability to swiftly adapt to dynamic conditions. Therefore, recent research endeavors have pivoted towards enhancing obstacle detection speed by employing CNN and particularly emphasizing the effectiveness of You Only Look Once (YOLO) [199]. YOLO is specifically designed for real-time object detection, and its variations have been employed to enhance success rates [200, 201, 202]. CNN offers promising capabilities in overcoming the limitations of traditional methods, showcasing greater adaptability and improved real-time performance in diverse and dynamic environments. However, given the restricted computational power available in small-scale cars, a balance must be achieved between accuracy and processing time. In [65], YOLOv3 and Tiny YOLO are selected as detection algorithms for the Go-CHART platform. [203] integrate GhostConv to the YOLOv4-tiny model to achieve faster detection, while in [204], YOLOv5 is used to detect cones and duckies within the Duckietown environment. Stereo image-based methods work by using a dual-camera system that captures images from slightly different angles, similar to the human binocular vision. This setup helps in perceiving depth and reconstructing three-dimensional scenes by analyzing the differences between the images from these cameras [205]. However, in the context of small-scale cars, constraints such as cost considerations and limited space availability restrict the usage of stereo cameras to select platforms, notably including RACECAR [18] and Autominy [23].
III-E2 LiDAR-based approach
@@
In addition to camera-based approaches, LiDAR is also widely used for obstacle avoidance systems. The applications of 2D LiDAR are prevalent in compact car platforms due to factors like size, price, weight, and overall compactness considerations. However, due to the existing computational limitations of these car platforms, innovative LiDAR data processing methods are often unavailable. Consequently, raw LiDAR point cloud data or minimally processed data is frequently used, involving procedures like imputing empty samples and cleansing noisy data, to serve as input for neural networks executing subsequent control tasks. In [48], the process of obstacle detection hinges on identifying the minimum value derived from a cluster of LiDAR points, following the determination of the position of the obstacle, a global path is generated and the vehicle is directed using MPC. After performing data cleaning, the minimum value obtained from the point cloud is adopted in [206] as the distance measurement to a potential collision object. Subsequently, a series of control command policies are executed based on this distance value. In [12], Follow the Gap (FTG) method [207] is employed for object detection and avoidance of an F1TENTH car. This technique involves computing the maximum gap present within the LiDAR point cloud and subsequently determining the steering control command. This command is derived from the heading angle directed toward the center of the largest gap to the orientation of the ego car, guiding the car toward the direction of the maximum gap center.
III-F Traffic sign recognition
@@
Traffic sign recognition can be seen as a special element of object detection tasks, requiring vehicles to initially detect and identify traffic signs before responding appropriately to a diverse array of road signs and signals. Unlike the generic object detection task, traffic sign recognition uniquely relies solely on cameras. This dependence stems from the necessity to detect and classify the visual characteristics of traffic signs. The process typically commences with dataset acquisition, wherein cameras capture traffic signs for later utilization. These captured images undergo a series of preprocessing steps to refine quality, eliminate noise, adjust lighting, and employ data augmentation techniques. Subsequently, ML models [208, 209, 210] are trained using these compiled datasets for detection and classification. However, research on traffic sign recognition is limited for small-scale cars due to the unavailability of standard datasets that may differ across platforms. Hence, researchers must prioritize dataset collection as an initial step to train the networks. Future research should intensify exploration in this direction, particularly focusing on mixing traffic sign recognition with other driving tasks like navigation, lane keeping, or path following.
III-G Racing

### v2

@@ IV-A3 Obstacle Recognition
Vision-based approaches
@@
The vision-based approach relies on images captured by cameras as the primary input for detection. This approach employs computer vision (CV) or ML algorithms, prominently leveraging CNN, for effective detection tasks. Within the realm of cameras used, the vision-based approach can be categorized into two subclasses: monocular and stereo. Monocular image-based methods rely on a single image for processing, while stereo methods utilize images captured by two synchronized cameras. Details about the camera sensors are discussed in Section III-A.
@@
Within monocular image-based methods, as outlined in [154], the obstacle detection task for small-scale cars is typically categorized into three primary domains: appearance-based, motion-based, and depth-based methods. With appearance-based methods, obstacles are typically identified as the foreground within images. The primary challenge lies in distinguishing relevant foreground or background elements based on established criteria, such as color discrepancies [155] or texture features [156]. In the work by [35], the RGB images captured by the monocular camera of a Duckiebot are initially transformed into the HSV colorspace. Subsequently, the color filter from OpenCV is employed to detect obstacles based on color differences. Its reliance on predefined color criteria significantly limits adaptability to diverse scenarios, such as environments with varying lighting or differently colored obstacles. This highlights the need for more robust methods that generalize well across varied settings, such as learning-based techniques. Motion-based obstacle detection refers to identifying obstacles or objects in an environment by analyzing motion vectors in the image, it involves comparing successive frames of images to determine alterations in position, velocity, or other motion-related characteristics. Objects that move or exhibit changes in their motion characteristics are then identified as potential obstacles. Optical flow analysis [157], background subtraction, differential methods. can be employed for motion-based obstacle detection [158, 159]. These methods excel in detecting moving obstacles but struggle with static hazards. Moreover, computational demands for real-time implementation can be prohibitive for small-scale platforms, calling for optimization techniques or lightweight motion estimation algorithms. Depth-based methods utilize depth information extracted from images to discern the distances and spatial arrangement of objects within the environment, aiding in accurate object detection [160, 161]. Employing conventional image processing techniques like this may fall short in meeting real-time application expectations, primarily due to their inability to swiftly adapt to dynamic conditions. Therefore, recent research endeavors have pivoted towards enhancing obstacle detection speed by employing CNN and particularly emphasizing the effectiveness of YOLO [92]. YOLO is specifically designed for real-time object detection, and its variations have been employed to enhance success rates [162, 163, 164]. CNN offers promising capabilities in overcoming the limitations of traditional methods, showcasing greater adaptability and improved real-time performance in diverse and dynamic environments. However, given the restricted computational power available in small-scale cars, a balance must be achieved between accuracy and processing time. In [62], YOLOv3 and Tiny YOLO are selected as detection algorithms for the Go-CHART platform. [165] integrate GhostConv to the YOLOv4-tiny model to achieve faster detection, while in [166], YOLOv5 is used to detect cones and duckies within the Duckietown environment. These aforementioned ML methods still face challenges in balancing detection accuracy and processing time on limited hardware. A potential solution involves further optimizing CNN architectures for edge deployment. Stereo image-based methods work by using a dual-camera system that captures images from slightly different angles, similar to the human binocular vision. This setup helps in perceiving depth and reconstructing three-dimensional scenes by analyzing the differences between the images from these cameras [167]. However, in the context of small-scale cars, constraints such as cost considerations and limited space availability restrict the usage of stereo cameras to select platforms, notably including RACECAR [24] and Autominy [27]. Exploring cost-effective stereo setups, such as single-camera depth estimation augmented with additional processing, could bridge this gap. Vision-based methods offer adaptability and versatility but face challenges related to environmental variability, computational overhead, and hardware limitations. Future work should explore lightweight CNN architectures, as highlighted in [168], where pruning techniques can significantly benefit computation-limited platforms. Additionally, hybrid depth estimation models and motion-aware algorithms hold promise for improving performance within the constraints of small-scale platforms.
LiDAR-based approaches
@@
In addition to camera-based approaches, LiDAR is also widely used for obstacle avoidance systems. The applications of 2D LiDAR are prevalent in compact car platforms due to factors like size, price, weight, and overall compactness considerations. However, due to the existing computational limitations of these car platforms, innovative LiDAR data processing methods are often unavailable. Consequently, raw LiDAR point cloud data or minimally processed data is frequently used, involving procedures like imputing empty samples and cleansing noisy data, to serve as input for neural networks executing subsequent control tasks. In [169] and [47], minimal processing of LiDAR data is employed, such as filtering and finding the nearest obstacle, to guide control policies. While computationally efficient, this approach provides limited contextual information, potentially hindering performance in complex scenarios with overlapping or clustered objects. Introducing advanced point cloud processing methods, such as voxel-based filtering or clustering, could enhance obstacle classification and avoidance capabilities. In [12], Follow the Gap (FTG) method [170] is employed for object detection and avoidance of an F1TENTH car. This technique involves computing the maximum gap present within the LiDAR point cloud and subsequently determining the steering control command. This approach is effective for reactive obstacle avoidance, especially in dynamic environments. However, it may struggle with narrow gaps. Combining FTG with predictive path planning or vision fusion could improve navigation in cluttered spaces. For small-scale cars, LiDAR remains a cornerstone of obstacle detection, offering reliable distance measurement and robustness. Enhancing LiDAR data processing with ML techniques, such as point cloud segmentation or recurrent neural networks, can address current limitations while maintaining real-time applicability.
Both vision-based and LiDAR-based approaches offer unique strengths but are subject to inherent trade-offs. Vision-based methods excel in rich environmental perception and are more cost-effective, but their susceptibility to environmental variability and computational demands can limit their reliability. LiDAR-based methods, while robust and precise, often lack the contextual depth provided by vision systems and can be hindered by computational simplicity. Integrating the two approaches into a multi-modal framework holds significant promise. For instance, vision can provide rich scene context, while LiDAR ensures precise spatial awareness. Techniques like sensor fusion and joint optimization of perception pipelines could leverage the complementary strengths of these modalities.
TABLE V: Comparison of methods used for obstacle detection.
Method Sensor Analysis Reference
Appearance-based Camera $\sqrt{}$ Simple, lightweight $\times$ Lighting or color variations [35]: Duckiebot
Motion-based Camera $\sqrt{}$ Effective for moving obstacles $\times$ High computational demand $\times$ Static obstacles [158, 159]: -
Depth-based Camera $\sqrt{}$ spatial awareness $\times$ High computational demand $\times$ Dynamic scenarios [161, 160]: -
CNN-based Camera $\sqrt{}$ High adaptability $\times$ Accuracy and speed trade-off [62]: Go-CHART; [165]: JetRacer; [166]: Duckiebot; [171]: Jetbot; [172, 173]: -
Minimum value LiDAR $\sqrt{}$ Efficient, robust for nearby objects $\times$ Limited contextual information [47, 169]: F1TENTH
FTG LiDAR $\sqrt{}$ Effective for reactive obstacle $\times$ Ineffective for narrow gaps [12]: F1TENTH
IV-A4 Traffic Sign Recognition
@@
Traffic sign recognition (TSR) is a specialized object detection task that enables autonomous vehicles to interpret road signage and adjust behavior accordingly. Unlike the generic object detection task, TSR uniquely relies solely on cameras, leveraging their capability to capture visual details. The process typically commences with dataset acquisition, wherein cameras capture traffic signs for later utilization. These captured images undergo a series of preprocessing steps to refine quality, eliminate noise, adjust lighting, and employ data augmentation techniques. Subsequently, ML models [174, 175, 93] are trained using these compiled datasets for detection and classification. However, research on TSR is limited for small-scale cars due to the unavailability of standard datasets that may differ across platforms. Hence, researchers must prioritize dataset collection as an initial step to train the networks. Future research should intensify exploration in this direction, particularly focusing on mixing TSR with other driving tasks like navigation, lane keeping, or path following, enabling vehicles to make informed, context-aware decisions in real time.
IV-B Planning
Path planning in autonomous driving refers to the process by which a self-driving vehicle determines a safe and efficient route from its current location to a desired destination while navigating through its environment [176]. This task involves creating a trajectory or path that the vehicle can follow, considering various factors such as obstacles, road conditions, and traffic regulations, and aims to ensure safe and reliable navigation while optimizing factors like travel time, and energy efficiency. Path planning algorithms vary in complexity and we specify them into traditional [177, 178, 179, 180] and ML-based techniques [181, 182].
Traditional methods
Traditional algorithms remain foundational in path planning due to their deterministic nature and efficiency in well-defined scenarios. For example, [183] applies Dijkstra’s algorithm [177] in conjunction with SLAM methods to extract global map information. The A* algorithm [178], noted for its optimality, is employed in [39] to plan optimal paths for racing cars, while [184] integrates Hybrid A* [185] with learned cost functions to address dynamic obstacles. However, reliance on precomputed cost functions may limit adaptability in novel environments, suggesting the need for adaptive cost-learning techniques. Genetic algorithms (GAs) offer advantages in unstructured environments, as explored in [186] and [187], where stochastic search capabilities generate global paths. However, high computational demands and convergence issues in real-time scenarios could be mitigated by hybrid approaches, combining GAs with deterministic methods like A*. To extend the planning capabilities beyond the sensor horizon, [188] utilizes a generative neural network trained on real-world data to predict occupancy maps beyond sensor limitations. This predictive capability assists in facilitating planning processes. Authors leverage the Rapidly-Exploring Random Tree (RRT) algorithm [179] to generate a global path, followed by a local planner that orchestrates trajectories until the end of the predicted map. While effective, traditional methods often struggle in dynamic or uncertain environments, emphasizing the importance of more adaptive frameworks.
ML-based methods
Path planning, treated as an optimization problem, has also seen the successful application of various ML methods, resulting in notably high-performance outcomes. For instance, [189] uses Q-Learning, a classical RL algorithm, to generate a global path for robots, which achieves a shorter and smoother path compared with the RRT algorithm, though challenges in scalability and training efficiency remain. Advanced reinforcement learning (RL) methods, such as Deep Q-Networks (DQN) and Policy Gradient algorithms, could enhance performance in future studies. A framework is introduced in [190] comprising a mapper, global policy, and local policy for image-based navigation for Jetbot. The mapper undergoes supervised learning using camera images to generate an occupancy grid map. Subsequently, a global policy employing DRL techniques takes both the map and the position as input to determine the long-term goal. Following this, a planner, employing the A* algorithm, computes the short-term goal, which is then forwarded to the local planner for further execution. Even with prior maps and acquired information, managing uncertainty during planning and driving remains a significant challenge. To address these uncertainties, [191] introduces a model-based RL algorithm incorporating a probabilistic dynamic model. This method aims to mitigate uncertainty during planning stages and prevent shortsighted decisions. In the related approach, [192] uses Bayesian Residual Policy Optimization (BRPO) with an ensemble of expert policies. Authors train the policy ensemble with BRPO to diminish overall system uncertainty, enabling safe navigation within partially observable environments. Demonstrated on the MuSHR platform, the proposed framework integrates a global localization system, ensuring destination-reaching capabilities while avoiding collisions with other moving vehicles.
For path planning, traditional methods excel in static and structured environments due to their computational efficiency and deterministic outputs. However, their inability to adapt to dynamic or uncertain conditions limits their application in complex real-world scenarios. Conversely, ML-based approaches offer flexibility and adaptability, capable of predicting and reacting to unforeseen changes, but they are computationally demanding and heavily reliant on large datasets. Future research should focus on hybrid frameworks that combine the computational efficiency of traditional methods with the adaptability of ML approaches. This could include transfer learning techniques to reduce data dependency, integration of predictive modeling to address uncertainties, and optimization of computational frameworks for real-time applications.
TABLE VI: Comparison of methods used for path planning.
Method Analysis Reference
Dijkstra’s Algorithm $\sqrt{}$ Deterministic $\sqrt{}$ Efficient in static environments $\times$ Inefficient for large graph [183]: -
A* $\sqrt{}$ Efficient in structured environments $\times$ Struggles in uncertain environment [39]: Deepracer
Hybrid A* $\sqrt{}$ Dynamic scenarios $\times$ Limited adaptability [184]: QCar
Genetic Algorithm $\sqrt{}$ High adaptability $\times$ High computational demand [186, 187]: -
RRT $\sqrt{}$ Efficient in high-dimensional spaces $\times$ Produce suboptimal paths [188]: RACECAR
DRL $\sqrt{}$ Handles dynamic environments $\times$ High computational demand [190]: Jetbot; [191]: MuSHR; [89]: Duckiebot; [189]: -
BRPO $\sqrt{}$ Robust in partially observable environments $\times$ High computational demand [192]: MuSHR
IL $\sqrt{}$ Efficient in known environments $\times$ Out-of-distribution error [193]: Donkeycar
IV-C Control
Once the path planning module establishes a desired path, the control module becomes the subsequent step. In this section, we will discuss the control methodologies that leverage the reference trajectory to compute control actions to navigate the car.
TABLE VII: Comparison of methods used for the control module.
Control Category Method Analysis Reference
Path Following Kinematic Controller $\sqrt{}$ Streamlines tuning efforts $\times$ Relies on accurate system modeling [194]: QCar; [195, 196]: -
PID Controller $\sqrt{}$ Simple, effective in structured systems $\times$ Limited adaptability [9]: Duckiebot
MPC $\sqrt{}$ Robust in structured environment $\times$ Requires maps $\times$ Unsuitable for dynamic environment [197]: MuSHR; [198, 199, 200]: -
IL $\sqrt{}$ Adaptable to varied scenarios $\times$ Struggles with unseen scenarios [201]: ART/ATK
DRL $\sqrt{}$ Adaptive to dynamic environment $\times$ High computational demand [202]: Autominy; [203]: JetRacer
Lane Keeping PID Controller $\sqrt{}$ Simple, effective in structured systems $\times$ Limited adaptability [9, 204, 152]: Duckiebot
Rule-based $\sqrt{}$ Straightforward, efficient $\times$ High computational demand [150]: Donkeycar
IL $\sqrt{}$ Adaptable to varied scenarios $\times$ Struggles with unseen scenarios [205, 206]: Duckiebot; [207]: -
DRL $\sqrt{}$ Handles dynamic environments $\times$ Limited adaptability [86, 85, 208]: Duckiebot; [42, 41]: Donkeycar
Car Following DRL $\sqrt{}$ Handles dynamic environments $\times$ Limited adaptability [86]: Duckiebot; [209]: Donkeycar; [210]: QCar
Overtaking MPC $\sqrt{}$ Robust in structured environment $\times$ Requires maps $\times$ Unsuitable for dynamic environment [211]: MuSHR
DRL $\sqrt{}$ Handles dynamic environments $\times$ Limited adaptability [85]: Duckiebot
Racing MPC $\sqrt{}$ Robust in structured environment $\times$ Requires maps $\times$ Unsuitable for dynamic environment [11, 54, 52]: Autorally; [20, 212, 213]: ORAC; [214]: F1TENTH
MPPI $\sqrt{}$ Real-time adaptation $\times$ High computational demand [50]: Autorally
Tube-MPC $\sqrt{}$ Safe and stable under dynamic conditions $\times$ High computational demand [215]: Autorally
LMPC $\sqrt{}$ Adaptability to uncertain dynamics $\sqrt{}$ Improvements with limited information $\times$ Large datasets $\times$ High computational demand [216, 146]: BARC
LPV-MPC $\sqrt{}$ Real-time implementation $\times$ Limited in highly nonlinear dynamics [121]: BARC
PAPC $\sqrt{}$ Improves safety $\times$ High computational demand [53]: Autorally
Pure Pursuit $\sqrt{}$ Lightweight, computationally efficient $\times$ Less robustness in high-speed [217]: TUNERCAR
Drifting LQR $\sqrt{}$ Effective in structured environment $\times$ Accurate dynamic models required [13, 218]: BARC
NMPC $\sqrt{}$ Improves control in dynamic environments $\times$ Computational cost [219]: BARC; [220]: RACECAR
Parking Rule-based $\sqrt{}$ Robust for well-defined cases $\times$ Limited flexibility [221, 222, 223]: -
Fuzzy $\sqrt{}$ Simple, interpretable $\times$ Adaptability in dynamic environments [224, 225, 226, 227, 228, 229, 230]: -
Neuro-fuzzy $\sqrt{}$ Adaptable to diverse scenarios $\times$ Training data required [231, 232]: -
GRBF $\sqrt{}$ Dynamic parking environments $\times$ Complex implementation [233]: -
IL $\sqrt{}$ Improves parking precision $\times$ High computational demand [234]: -
DRL $\sqrt{}$ Highly adaptable to dynamic environments $\times$ High computational demand [235, 236]: -
Cooperative driving Centralized Control $\sqrt{}$ Globally optimized coordination $\sqrt{}$ Effective for requiring high precision $\times$ High computational demand $\times$ Limited scalability [28]: CoRoLa; [40]: DeepRacer; [58]: UDSSC; [237]: Jetbot; [238]: QCar
Decentralized Control $\sqrt{}$ Robustness and scalability $\times$ Effective communication protocols required $\times$ Potential for suboptimal solutions [63]: CamMini; [239]: QCar; [240]: MuSHR; [241]: $\mu$Car; [56, 242, 243, 21]: UDSSC; [244]: Jetbot
IV-C1 Path Following
Path following is the most simple control task for small-scale cars, it necessitates a continual adjustment of the movements based on real-time sensor feedback and environmental alterations, ensuring the vehicle maintains the desired trajectory accurately. Solutions for path following encompass a wide spectrum, including classical control methods like the PID controller, kinematic controller [195, 196, 194], MPC [198, 199, 200, 197], and extending to ML methods [201, 202, 190, 203].
In [194], an adaptive trajectory tracking control scheme is introduced with adjustable gains to facilitate adherence to predefined paths. The designed adaptive control gains aimed to streamline tuning efforts, augment the convergence rate of tracking errors, and elevate trajectory tracing performance. However, a key disadvantage is its reliance on accurate system modeling and gain adjustment, which may struggle under significant environmental disturbances or non-linear dynamics. The strategy introduced by [197] involves expert interventions with pre-existing LiDAR-derived maps to further improve performance. Nevertheless, the dependency on pre-existing maps and the requirement for expert input make it less suitable for dynamic or unstructured environments. For ML-based methods, in the work presented by [201], error states relative to reference trajectories are determined using vehicle state data obtained from sensor fusion involving IMU and GPS with an EKF. A Neural Network (NN) is trained using IL methods with a dataset collected from human-controlled and MPC-controlled driving scenarios. Trained NN maps error states to control commands. While this method significantly reduces the need for complex hand-crafted control laws, it inherits biases from the training data, especially when collected from suboptimal human drivers. To mitigate the need for extensive datasets, [202] applies Deep Deterministic Policy Gradient (DDPG) algorithm to guide an Autominy car along the desired path, aiming to minimize cross-track errors. Similarly, in the study by [190], a global policy is determined using DRL, while a local planner processes received images with another DRL agent to derive the final action for the path following. To address the data inefficiency challenges with RL, [203] then introduces a policy gradient-based policy optimization framework leveraging a first-principles model for path following task with a JetRacer. This framework facilitated the learning of precise control policies with limited real-world data. While promising, the reliance on first-principles modeling may limit its adaptability to systems with uncertain or highly non-linear dynamics. Furthermore, the approach assumes access to accurate and reliable simulation environments, which might not always reflect real-world variability.
IV-C2 Lane Keeping
When small-scale cars drive on tracks without obstacles, the optimal planned trajectory typically aligns with the center of the right lane. Together with the lane detection methods discussed in Section IV-A2, the primary objective is to guide the vehicle within the designated lane, adhering to predefined tolerances for deviations from the lane center. This task requires the car to dynamically adjust its path in real time, based on its position relative to the lane. Consequently, the task heavily depends on visual input, typically utilizing RGB cameras to monitor the lane position of the vehicle continuously. This process, commonly referred to as lane keeping for small-scale vehicles, has been a central focus of extensive research across various platforms [9, 10, 23]. Research in lane keeping can be broadly categorized into two approaches: traditional methods and ML-based methods.
Traditional methods
Traditional control methods regulate the trajectory using outputs from the perception module, ensuring it remains within the designated lane. Algorithms such as PID controllers are commonly used to adjust the steering angle based on deviations from the lane center. For instance, in [9], the PID controller utilizes lateral displacement and angular offset estimates from the perception pipeline to minimize deviations in real time. Similarly, [150] integrates a rule-based approach to compute steering angles directly from detected lane lines. These methods are straightforward and interpretable, making them suitable for simple lane keeping tasks. However, their reliance on precise input metrics and static tuning parameters limits their effectiveness in dynamic environments or when system dynamics are nonlinear.
ML-based methods
ML-based control methods enhance lane keeping by either augmenting traditional control systems or replacing them with more adaptive algorithms. For example, DRL has been successfully used to replace traditional controllers. [86, 85] employ DRL algorithms to process outputs from the perception module of [9], achieving superior performance in multitask driving scenarios beyond lane keeping. Additionally, these systems improve Sim2Real transfer capabilities, as demonstrated by [42], where a modular system integrates a list of DRL algorithms, utilizing these latent features from the perception module as inputs to control the Donkeycar. While ML-based control methods provide adaptability and can learn optimal policies from data, they are often sensitive to hyperparameter tuning and may require extensive training.
IV-C3 Car Following
In addition to navigating tracks without other vehicles, a key driving scenario for autonomous vehicles involves operating alongside other vehicles on the road. One crucial capability in such scenarios is car following, which refers to the ability to maintain a safe and appropriate distance from the vehicle ahead. This requires tracking the movements of the leading vehicle and dynamically adjusting speed and position to ensure safe, comfortable, and efficient driving. The literature on car following maneuvers with small-scale vehicles remains relatively limited, though several notable studies have emerged. A DDPG algorithm with an extended look-ahead approach is proposed in [209] for longitudinal and lateral control within vehicle platooning scenarios. For perception, LiDAR technology gauges inter-vehicle distance, while an IMU provides acceleration data. Additionally, a V2V system employing Wi-Fi communication transmits leader information to the follower vehicle. Similarly, in [210], a Cooperative Adaptive Cruise Control (CACC) system is implemented utilizing Deep Q-learning (DQN) [245]. This system enables the follower vehicle to dynamically adapt and maintain appropriate inter-vehicular distances. Notably, both these studies rely on V2V communication to achieve effective car following behavior, which increases system complexity and costs. In contrast, [86] introduces a new perspective by realizing multitasking driving, encompassing car following and lane keeping, solely relying on visual sensors without the need for communication with other vehicles. In this approach, a pattern of circles is affixed behind the leading Duckiebot, aiding the ego Duckiebot in sensing distance and speed to mimic human driver behavior. A modular system is deployed, where the perception module extracts compact affordance information, and a DRL algorithm controls the Duckiebot based on this information. Despite its simplicity, this system yields promising results.
IV-C4 Overtaking
Overtaking, like car following, is another critical driving behavior that involves interaction with other vehicles on the track. It refers to the maneuver where an autonomous vehicle changes lanes or positions itself to safely pass another vehicle traveling in the same direction on the road. The purpose of overtaking is to move ahead of the slower vehicle safely and efficiently while maintaining proper traffic flow. For small-scale cars, overtaking is particularly challenging due to the need for precise localization to avoid collisions throughout the maneuver. Consequently, limited research has been conducted on this task. A first contribution comes from [85], where authors employ a modular system equipped solely with onboard sensors to execute diverse driving tasks, encompassing lane keeping and overtaking for Duckiebots. The framework leverages traditional machine vision technologies to acquire compact affordance information, as discussed in [9], and utilizes LiDAR sensors for distance estimation relative to other vehicles. Afterward, the Long Short-Term Memory Soft Actor-Critic (LSTM-SAC) algorithm assumes the role of the controller. Notably, LSTM aids the agent in recognizing distinct phases within the overtaking process. Demonstrated results affirm the efficacy of this proposed framework, managing both lane keeping and overtaking tasks, and showcasing superior performance when benchmarked against baselines. Another work regarding the overtaking task is [211], where a dual control approach with MPC towards active uncertainty reduction is proposed. This approach automatically balances the exploration-exploitation trade-off, enabling the MuSHR car to actively minimize uncertainty concerning the hidden states of other agents without compromising expected planning performance. The difference compared with [85] from the hardware side is that the vehicle being overtaken will yield, and the usage of a known grid map of the track. While these methods showcase promising results, they highlight the current research limitations in overtaking tasks for small-scale cars. Future work should prioritize the development of more generalizable frameworks capable of handling diverse overtaking scenarios without relying on assumptions such as yielding vehicles or predefined maps.
IV-C5 Racing